# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 3.1.3 Reinforcement Learning
Single-Turn RL Reinforcement learning from human feedback (RLHF) aligns LLMs with human preferences by using reward models trained on human-generated rankings, significantly improving response quality and safety. For instance, a 1.3B-parameter InstructGPT fine-tuned via RLHF outperformed a 175B-parameter GPT-3 in human evaluations, exhibiting greater truthfulness and reduced toxicity [190]. Reinforcement learning from AI feedback (RLAIF) extends RLHF by employing AI-generated feedback signals instead of human labels. Anthropic’s Constitutional AI [240] notably trains models using AI-authored principles and self-critique, successfully reducing harmful behaviors without direct human supervision. Both RLHF and RLAIF have proven effective for aligning single-turn LLM responses with human values.
@@
Early RLHF successes largely utilized Proximal Policy Optimization (PPO), an on-policy method offering stable and sample-efficient fine-tuning of large models. For example, OpenAI’s InstructGPT [190] employed PPO to optimize human-preference-based reward models, significantly enhancing helpfulness and reducing harmful outputs with minimal performance loss on traditional NLP benchmarks.
@@
While PPO-based RLHF is effective, it can be complex due to reward modeling and tuning. Direct Preference Optimization (DPO) [191] simplifies this by converting RLHF into a supervised classification task, analytically deriving optimal policies from preference models. Context-aware Preference Optimization (CaPO, [24]) extends DPO by integrating context-aware preferences, promoting correct context utilization. Additionally, Action-Based Contrastive Self-Training (ACT) [192] applies DPO quasi-online, enhancing LLMs’ clarifying behaviors in ambiguous contexts while preserving DPO’s simplicity.
@@
Many previous studies have investigated RLHF methods for enhancing multi-turn interactions in LLMs [24, 30, 23, 28, 46, 20, 22, 65, 109, 122, 127, 125, 150]. By horizontally comparing these works, however, a nuanced and somewhat contradictory picture emerges regarding RLHF’s effectiveness. On one hand, specialized domains such as medical QA [122, 127] and coding tasks [22, 65] demonstrate substantial performance gains from carefully implemented RLHF approaches like DPO and interactive feedback loops. Similarly, the Parrot framework [24] underscores how explicitly training on negative examples can guide models towards better contextual alignment and fewer repetitive errors.
@@
On the other hand, broader analyses from MINT [23] and MT-Bench 101 [30] reveal that generalized RLHF methods frequently yield limited or even negative impacts in multi-turn scenarios. This discrepancy highlights the inherent limitations of RLHF methods in terms of generalizability, emphasizing the necessity for tailored RL strategies explicitly designed for multi-turn interactions.
@@
Multi-Turn RL Recent research has therefore extended preference optimization and reinforcement learning techniques to optimize entire conversation trajectories rather than single responses. Below, we review key developments organized by technique.
@@
A straightforward way to align multi-turn behavior is to generalize the single-turn preference optimization objective across an entire dialogue trajectory. Direct Multi-Turn Preference Optimization (DMPO) [193] is one such approach that adapts DPO to multi-turn agent tasks. A technical obstacle in multi-turn DPO is that the partition function (normalization factor) no longer cancels out between preferred and dispreferred trajectory pairs, complicating the loss computation. DMPO addresses this by re-formulating the optimization: it replaces the per-response probability ratio with a state-action occupancy measure and normalizes for sequence length differences. The result is a novel DMPO loss that comes with theoretical justification for multi-turn settings.
Related work by Xiong et al., (2024) [52], which we covered in §2.1.2, introduced an iterative multi-turn preference learning framework with two implementations: Multi-turn DPO (M-DPO) and Multi-turn KTO (M-KTO). (Kahneman-Tversky Optimization (KTO) is a recently proposed RLHF approach inspired by prospect theory, designed to better align model outputs with human preferences by explicitly modeling decision-making biases [241].) Their focus was on mathematical reasoning agents that utilize tools (code interpreters) across multiple turns. Because single-turn preference methods did not fully capture multi-step reasoning quality, they extended DPO and the prospect-theoretic KTO loss to handle entire solution trajectories. In this framework, a reward model (augmented with tool feedback) judges the quality of a full reasoning trace, and the model is tuned to prefer better traces. Both M-DPO and M-KTO yielded substantial performance gains on math problem benchmarks: for instance, a 7B model’s accuracy on GSM8K math questions jumped from 77.5% to 83.9% after multi-turn preference optimization.
@@
Multi-turn interactions intensify the credit assignment problem—attributing outcomes to specific decisions made across several dialogue turns. Hierarchical reinforcement learning addresses this challenge by structuring decision-making across multiple abstraction levels. ArCHer [194] employs a two-level architecture: a high-level module managing dialogue-turn granularity and a low-level module generating tokens within each turn. Its high-level off-policy learner computes Q-values based on accumulated dialogue rewards, guiding low-level updates via actor-critic methods. ArCHer achieves substantial sample efficiency, outperforming non-hierarchical baselines by approximately 100× in terms of training samples.
DeepMind’s SCoRe [195], an extension of ArCHer, explicitly targets self-correction by training models on synthetic dialogues to recognize and rectify errors across two-turn sequences. SCoRe significantly advances benchmarks for self-correction in math and coding tasks. Overall, hierarchical frameworks like ArCHer and SCoRe effectively address multi-turn credit assignment through structured training and specialized rewards.
@@
Another promising direction frames multi-turn LLM alignment as an off-policy value-learning problem. REFUEL [196] addresses covariate shift issues in multi-turn RLHF by iteratively training a Q-value function on accumulated trajectories, directly regressing future cumulative rewards. Unlike standard on-policy methods, REFUEL continuously leverages self-generated dialogues, ensuring accurate value estimation for inference-time states. Empirically, REFUEL outperformed state-of-the-art methods like DPO on long dialogue benchmarks; notably, an 8B-parameter REFUEL model surpassed a 70B-parameter model fine-tuned via single-turn methods, highlighting the value of explicit long-term reward modeling for improved conversational coherence.
@@
Effectively assessing multi-turn RL algorithms requires specialized benchmarks that capture long-term interaction and credit assignment. LMRL-Gym [242] introduces a suite of interactive language tasks—ranging from open-ended dialogues to text-based games—to measure critical skills such as intentionality, information-gathering, and strategic planning over extended interactions. By providing both offline and online RL evaluation frameworks, LMRL-Gym enables systematic benchmarking of multi-turn RL improvements. Similarly, SWEET-RL [243] introduces ColBench, a set of collaborative human-AI tasks emphasizing multi-turn dialogue and reasoning, where traditional RL struggles due to delayed, sparse feedback. SWEET-RL addresses this by training a turn-wise advantage critic using additional training-time signals, significantly improving performance on collaborative tasks such as coding and UI design conversations. Together, LMRL-Gym and SWEET-RL highlight the importance of tailored benchmarks and evaluation methods to reliably advance LLMs’ multi-turn conversational abilities.
@@ 3.1.4 New Architectures
Some researchers have questioned whether inherent limitations of the Transformer architecture itself might be responsible for observed performance degradation in complex, multi-turn scenarios [244]. Motivated by this concern, in addition to advances in contextual learning, supervised fine-tuning, and reinforcement learning, recent efforts have explored optimizing LLM architectures to specifically enhance performance in multi-turn interactions.
@@
Cached Transformers [197] introduce a novel model architecture that extends the traditional Transformer by incorporating a Gated Recurrent Cache (GRC) into its self-attention mechanism. This differentiable memory cache compresses historical token representations into fixed-length vectors that are continuously updated through gating, enabling the model to attend efficiently to both past and current tokens. By effectively capturing long-range dependencies without significant computational overhead, Cached Transformers improve performance on various language tasks. In multi-turn conversations, this mechanism can help LLMs maintain coherent and contextually rich dialogue histories by providing a persistent, efficient memory of earlier interactions, thereby enhancing the model’s ability to reference and build upon past conversational turns.
@@
Beyond caching, researchers are exploring stateful transformer designs that maintain an internal dialogue state. For example, Wu et al. (2023) [198] introduced a memory-augmented transformer (MemBART) that carries a “memory state” alongside the normal model hidden state, updated at each turn. MemBART employs a dual attention stream to separately handle memory reading and writing, along with a residual gated update mechanism that determines how much past information to retain versus update at each timestep. This design allows the model to efficiently store and retrieve important historical context without needing excessively large input windows, thereby enhancing its ability to maintain coherent multi-turn conversations with lower computational overhead and improved latency.
@@
Recent advances in long-context language modeling have also leveraged recurrence to extend Transformers’ effective context. Transformer-XL [199] reuses hidden states from previous segments with a novel relative positional encoding to mitigate context fragmentation. The Recurrent Memory Transformer [200] augments this idea by inserting dedicated memory tokens into the sequence, which are recurrently updated to store global information. He et al. (2024) [201] presents a transformer framework that mimics the brain’s memory hierarchy by segmenting information into sensory, short-term, and long-term layers, thereby facilitating the processing of lengthy contexts with lower computational overhead. Similarly, RWKV [202] reformulates attention in an RNN-like manner, achieving linear complexity while retaining Transformer parallelism. Enhancing RWKV-based models [203], recent work introduces adaptive gating and position-aware convolutional shifts to dynamically regulate inter-token information flow. All four approaches share the common goal of overcoming fixed-length limitations by propagating information across segments, thereby capturing long-term dependencies more efficiently while balancing training parallelism with inference efficiency.
@@ 3.2.1 Memory-Augmented Methods
MemPrompt [204] demonstrates an early external memory approach where the system records pairs of misunderstood inputs and corresponding user corrections in a dynamic memory bank. When a similar query arises later, the stored corrective feedback is retrieved and appended to the prompt, guiding the model toward a more accurate interpretation. In a related effort, Wu et al. (2024) [205] proposes a unified framework that decomposes memory design into indexing, retrieval, and reading stages. Their work introduces detailed optimizations—including session decomposition, fact-augmented key expansion, and time-aware query expansion—that significantly enhance memory recall and downstream question-answering accuracy, even for models designed with extended contexts.
Taking inspiration from human cognition, Pink et al. (2025) [206] argues for the integration of episodic memory into LLMs. Their framework emphasizes long-term storage, explicit reasoning, and instance-specific detail capture, outlining four research directions: discretizing continuous interactions into episodes, retrieving relevant past experiences, consolidating episodic traces into generalized knowledge, and establishing benchmarks for evaluation. Meanwhile, hierarchical memory structures offer another avenue for improvement. Rezazadeh et al. (2024) [207] introduces MemTree, a dynamic tree-based system that aggregates dialogue content into hierarchical nodes to enable efficient retrieval and improved long-term reasoning.
@@ 3.2.2 Retrieval Augmented Generated
Retrieval-augmented generation (RAG) is an advanced framework introduced by Lewis et al. (2020)  [245] that enhances the capabilities of NLP models by integrating external knowledge sources into the generation process. The RAG integration allows LLMs to access up-to-date and domain-specific information, improving the accuracy and relevance of their responses. It also helps mitigate AI hallucinations by grounding outputs in reliable data.
@@
Studies have applied retrieval-based architectures to multi-turn dialogue systems, enabling the generation of more informative and factual responses by conditioning on both the user’s input and relevant external documents. For instance, Wizard of Wikipedia [208] incorporated retrieval into each dialogue turn, resulting in significantly higher factual accuracy and user engagement compared to non-retrieval baselines. Similarly, Komeili et al. (2021) [209] explicitly generates search queries from the dialogue context to pull in up-to-date knowledge (e.g. via internet search) and then conditions the response on the retrieved results. Other variants, like [210], retrieve from a fixed knowledge base or enterprise documents, using dense vector search to find passages related to the user’s query, where the generator model then conditions on both the dialogue context and the fetched evidence to produce a response. BlenderBot 2.0 [211] extended the RAG idea by integrating both internet search and a long-term memory component, allowing the model to sustain coherent conversations across multiple turns and sessions while retrieving past facts when needed. This allows the system to handle context dependencies that go beyond the immediate dialogue window. Overall, RAG offers a principled mechanism for overcoming the context length and memory limitations of LLMs, making it a valuable technique for improving dialogue coherence, answer accuracy, and user trust in multi-turn interactions.
@@
To assess the effectiveness of RAG systems in multi-turn conversational settings, benchmarks such as MTRAG [246], CORAL [247], and RAD-Bench [248] have been developed. MTRAG comprises 110 conversations averaging 7.7 turns each, totaling 842 tasks across four domains and incorporates diverse question types (factoid, comparison, explanation, etc.), varying answerability (answerable, partially answerable, unanswerable), and multi-turn dynamics (follow-up and clarification questions), providing a comprehensive evaluation framework for multi-turn RAG systems. The benchmark emphasizes active retrieval, where relevant documents are dynamically fetched based on user inquiries throughout the conversation, simulating a more realistic conversational experience. Similarly, CORAL offers a large-scale benchmark designed to assess RAG systems in realistic multi-turn conversational settings, supporting tasks such as passage retrieval, response generation, and citation labeling. RAD-Bench provides a frame to assess multi-turn LLMs’ capabilities in augmented generation with retrieved context in multi-turn scenarios with both Retrieval Synthesis and Retrieval Reasoning mechanisms. The evaluation framework contains 89 multi-turn question samples sampled across six practical scenarios inspired by human-LLM multi-turn dialogue interactions requiring retrieved context to complete tasks.

### v5

@@ 3.1.3 Reinforcement Learning
Reinforcement learning from human feedback (RLHF) aligns LLMs with human preferences via reward models trained on human rankings, substantially improving response quality and safety. Notably, a 1.3B-parameter InstructGPT fine-tuned with RLHF outperformed a 175B-parameter GPT-3 in human evaluations, exhibiting greater truthfulness and reduced toxicity (Ouyang et al., 2022). Reinforcement learning from AI feedback (RLAIF) extends this paradigm by substituting AI-generated feedback for human labels; Anthropic’s Constitutional AI (Bai et al., 2022) exemplifies this approach, training models via AI-authored principles and self-critique to reduce harmful behaviors without direct human supervision. Both RLHF and RLAIF have proven effective for aligning single-turn LLM responses with human values.
@@
Early RLHF successes largely utilized Proximal Policy Optimization (PPO), an on-policy method offering stable and sample-efficient fine-tuning of large models. For example, OpenAI’s InstructGPT (Ouyang et al., 2022) employed PPO to optimize human-preference-based reward models, significantly enhancing helpfulness and reducing harmful outputs with minimal performance loss on traditional NLP benchmarks.
@@
While PPO-based RLHF is effective, it can be complex due to reward modeling and tuning. Direct Preference Optimization (DPO) (Rafailov et al., 2023) simplifies this by converting RLHF into a supervised classification task, analytically deriving optimal policies from preference models. Context-aware Preference Optimization (CaPO, (Sun et al., 2024)) extends DPO by integrating context-aware preferences, promoting correct context utilization. Additionally, Action-Based Contrastive Self-Training (ACT) (Chen et al., 2025a) applies DPO quasi-online, enhancing LLMs’ clarifying behaviors in ambiguous contexts while preserving DPO’s simplicity.
@@
Many prior studies have investigated RLHF for enhancing multi-turn LLM interactions (Sun et al., 2024; Bai et al., 2024; Wang et al., 2024f; Lù et al., 2024; Du et al., 2025; Wu et al., 2024b; Yang et al., 2023; Chen et al., 2025d; Bao et al., 2023; Yang et al., 2024; Ye et al., 2023; Chen et al., 2024f; Scarlatos et al., 2024). A horizontal comparison of these works, however, reveals a nuanced and somewhat contradictory picture. On one hand, specialized domains such as medical QA (Yang et al., 2024; Ye et al., 2023) and coding (Yang et al., 2023; Chen et al., 2025d) demonstrate substantial gains from carefully implemented approaches like DPO and interactive feedback loops. Similarly, the Parrot framework (Sun et al., 2024) highlights how explicitly training on negative examples can steer models toward better contextual alignment and fewer repetitive errors.
@@
On the other hand, broader analyses from MINT (Wang et al., 2024f) and MT-Bench 101 (Bai et al., 2024) reveal that generalized RLHF methods frequently yield limited or even negative impacts in multi-turn scenarios. This discrepancy highlights the inherent limitations of RLHF methods in terms of generalizability, emphasizing the necessity for tailored RL strategies explicitly designed for multi-turn interactions.
Multi-Turn RL
@@
Recent research has therefore extended preference optimization and reinforcement learning techniques to optimize entire conversation trajectories rather than single responses. Below, we review key developments organized by technique.
@@
A straightforward way to align multi-turn behavior is to generalize the single-turn preference optimization objective across an entire dialogue trajectory. Direct Multi-Turn Preference Optimization (DMPO) (Shi et al., 2024b) is one such approach that adapts DPO to multi-turn agent tasks. A technical obstacle in multi-turn DPO is that the partition function (normalization factor) no longer cancels out between preferred and dispreferred trajectory pairs, complicating the loss computation. DMPO addresses this by re-formulating the optimization: it replaces the per-response probability ratio with a state-action occupancy measure and normalizes for sequence length differences. The result is a novel DMPO loss that comes with theoretical justification for multi-turn settings.
Related work by Xiong et al. (2025), which we covered in §2.1.2, introduced an iterative multi-turn preference learning framework with two implementations: Multi-turn DPO (M-DPO) and Multi-turn KTO (M-KTO). (Kahneman-Tversky Optimization (KTO) is a recently proposed RLHF approach inspired by prospect theory, designed to better align model outputs with human preferences by explicitly modeling decision-making biases (Ethayarajh et al., 2024).) Their focus was on mathematical reasoning agents that utilize tools (code interpreters) across multiple turns. Because single-turn preference methods did not fully capture multi-step reasoning quality, they extended DPO and the prospect-theoretic KTO loss to handle entire solution trajectories. In this framework, a reward model (augmented with tool feedback) judges the quality of a full reasoning trace, and the model is tuned to prefer better traces. Both M-DPO and M-KTO yielded substantial performance gains on math problem benchmarks: for instance, a 7B model’s accuracy on GSM8K math questions jumped from 77.5% to 83.9% after multi-turn preference optimization.
@@
Multi-turn interactions intensify the credit assignment problem—attributing outcomes to specific decisions made across several dialogue turns. Hierarchical reinforcement learning addresses this challenge by structuring decision-making across multiple abstraction levels. ArCHer (Zhou et al., 2024b) employs a two-level architecture: a high-level module managing dialogue-turn granularity and a low-level module generating tokens within each turn. Its high-level off-policy learner computes Q-values based on accumulated dialogue rewards, guiding low-level updates via actor-critic methods. ArCHer achieves substantial sample efficiency, outperforming non-hierarchical baselines by approximately 100× in terms of training samples.
DeepMind’s SCoRe (Kumar et al., 2025), an extension of ArCHer, explicitly targets self-correction by training models on synthetic two-turn dialogues to recognize and rectify errors. SCoRe yields significant gains on self-correction benchmarks in math and coding. Together, hierarchical frameworks like ArCHer and SCoRe effectively address multi-turn credit assignment through structured training and specialized reward design.
@@
Another promising direction frames multi-turn LLM alignment as an off-policy value-learning problem. REFUEL (Gao et al., 2025b) addresses covariate shift issues in multi-turn RLHF by iteratively training a Q-value function on accumulated trajectories, directly regressing future cumulative rewards. Unlike standard on-policy methods, REFUEL continuously leverages self-generated dialogues, ensuring accurate value estimation for inference-time states. Empirically, REFUEL outperformed state-of-the-art methods like DPO on long dialogue benchmarks; notably, an 8B-parameter REFUEL model surpassed a 70B-parameter model fine-tuned via single-turn methods, highlighting the value of explicit long-term reward modeling for improved conversational coherence.
@@
Effectively evaluating multi-turn RL algorithms requires specialized benchmarks that capture long-horizon interaction and credit assignment. LMRL-Gym (Abdulhai et al., 2025) provides a suite of interactive language tasks—spanning open-ended dialogues to text-based games—designed to measure intentionality, information-gathering, and strategic planning over extended interactions, supporting both offline and online RL evaluation. Similarly, SWEET-RL (Zhou et al., 2025) introduces ColBench, a set of collaborative human-AI tasks emphasizing multi-turn dialogue and reasoning, where sparse delayed feedback makes traditional RL unreliable. SWEET-RL addresses this by training a turn-wise advantage critic with additional training-time signals, yielding significant gains on collaborative tasks such as coding and UI design. Together, LMRL-Gym and SWEET-RL underscore the importance of tailored benchmarks and evaluation protocols for reliably advancing multi-turn conversational abilities.
@@ 3.1.4 New Architectures
Some researchers have questioned whether inherent limitations of the Transformer architecture itself might be responsible for observed performance degradation in complex, multi-turn scenarios (Dziri et al., 2023). Motivated by this concern, in addition to advances in contextual learning, supervised fine-tuning, and reinforcement learning, recent efforts have explored optimizing LLM architectures to specifically enhance performance in multi-turn interactions.
@@
Cached Transformers (Zhang et al., 2024d) introduce a novel model architecture that extends the traditional Transformer by incorporating a Gated Recurrent Cache (GRC) into its self-attention mechanism. This differentiable memory cache compresses historical token representations into fixed-length vectors that are continuously updated through gating, enabling the model to attend efficiently to both past and current tokens. By effectively capturing long-range dependencies without significant computational overhead, Cached Transformers improve performance on various language tasks. In multi-turn conversations, this mechanism can help LLMs maintain coherent and contextually rich dialogue histories by providing a persistent, efficient memory of earlier interactions, thereby enhancing the model’s ability to reference and build upon past conversational turns.
@@
Beyond caching, researchers are exploring stateful transformer designs that maintain an internal dialogue state. For example, Wu and Yu (2024) introduce a memory-augmented transformer (MemBART) that carries a “memory state” alongside the normal model hidden state, updated at each turn. MemBART employs a dual attention stream to separately handle memory reading and writing, along with a residual gated update mechanism that determines how much past information to retain versus update at each timestep. This design allows the model to efficiently store and retrieve important historical context without needing excessively large input windows, thereby enhancing its ability to maintain coherent multi-turn conversations with lower computational overhead and improved latency.
@@
Recent advances in long-context language modeling have also leveraged recurrence to extend Transformers’ effective context. Transformer‑XL (Dai et al., 2019) reuses hidden states from previous segments with a novel relative positional encoding to mitigate context fragmentation. The Recurrent Memory Transformer (Bulatov et al., 2022) augments this idea by inserting dedicated memory tokens into the sequence, which are recurrently updated to store global information. He et al. (2025) present a transformer framework that mimics the brain’s memory hierarchy by segmenting information into sensory, short-term, and long-term layers, thereby facilitating the processing of lengthy contexts with lower computational overhead. Similarly, RWKV (Peng et al., 2023) reformulates attention in an RNN-like manner, achieving linear complexity while retaining Transformer parallelism. Enhancing RWKV-based models (Pan, 2025), recent work introduces adaptive gating and position-aware convolutional shifts to dynamically regulate inter-token information flow. All four approaches share the common goal of overcoming fixed-length limitations by propagating information across segments, thereby capturing long-term dependencies more efficiently while balancing training parallelism with inference efficiency.
@@ 3.2.1 Memory-Augmented Methods
MemPrompt (Madaan et al., 2022) demonstrates an early external memory approach where the system records pairs of misunderstood inputs and corresponding user corrections in a dynamic memory bank. When a similar query arises later, the stored corrective feedback is retrieved and appended to the prompt, guiding the model toward a more accurate interpretation. In a related effort, Wu et al. (2025) propose a unified framework that decomposes memory design into indexing, retrieval, and reading stages. Their work introduces detailed optimizations—including session decomposition, fact-augmented key expansion, and time-aware query expansion—that significantly enhance memory recall and downstream question-answering accuracy, even for models designed with extended contexts.
Taking inspiration from human cognition, Pink et al. (2025) argue for the integration of episodic memory into LLMs. Their framework emphasizes long-term storage, explicit reasoning, and instance-specific detail capture, outlining four research directions: discretizing continuous interactions into episodes, retrieving relevant past experiences, consolidating episodic traces into generalized knowledge, and establishing benchmarks for evaluation. Meanwhile, hierarchical memory structures offer another avenue for improvement. Rezazadeh et al. (2025) introduce MemTree, a dynamic tree-based system that aggregates dialogue content into hierarchical nodes to enable efficient retrieval and improved long-term reasoning.
@@ 3.2.2 Retrieval Augmented Generated
Retrieval-augmented generation (RAG) is an advanced framework introduced by Lewis et al. (2020) that enhances the capabilities of NLP models by integrating external knowledge sources into the generation process. The RAG integration allows LLMs to access up-to-date and domain-specific information, improving the accuracy and relevance of their responses. It also helps mitigate AI hallucinations by grounding outputs in reliable data.
@@
Studies have applied retrieval-based architectures to multi-turn dialogue systems, enabling the generation of more informative and factual responses by conditioning on both the user’s input and relevant external documents. For instance, Wizard of Wikipedia (Dinan et al., 2018) incorporated retrieval into each dialogue turn, resulting in significantly higher factual accuracy and user engagement compared to non-retrieval baselines. Similarly, Komeili et al. (2022) explicitly generate search queries from the dialogue context to pull in up-to-date knowledge (e.g. via internet search) and then conditions the response on the retrieved results. Other variants, like Karpukhin et al. (2020), retrieve from a fixed knowledge base or enterprise documents, using dense vector search to find passages related to the user’s query, where the generator model then conditions on both the dialogue context and the fetched evidence to produce a response. BlenderBot 2.0 (Xu et al., 2022) extend the RAG idea by integrating both internet search and a long-term memory component, allowing the model to sustain coherent conversations across multiple turns and sessions while retrieving past facts when needed. This allows the system to handle context dependencies that go beyond the immediate dialogue window. Overall, RAG offers a principled mechanism for overcoming the context length and memory limitations of LLMs, making it a valuable technique for improving dialogue coherence, answer accuracy, and user trust in multi-turn interactions.
@@
To assess the effectiveness of RAG systems in multi-turn conversational settings, benchmarks such as MTRAG (Katsis et al., 2025), CORAL (Cheng et al., 2025), and RAD-Bench (Kuo et al., 2025) have been developed. MTRAG comprises 110 conversations averaging 7.7 turns each, totaling 842 tasks across four domains and incorporates diverse question types (factoid, comparison, explanation, etc.), varying answerability (answerable, partially answerable, unanswerable), and multi-turn dynamics (follow-up and clarification questions), providing a comprehensive evaluation framework for multi-turn RAG systems. The benchmark emphasizes active retrieval, where relevant documents are dynamically fetched based on user inquiries throughout the conversation, simulating a more realistic conversational experience. Similarly, CORAL offers a large-scale benchmark designed to assess RAG systems in realistic multi-turn conversational settings, supporting tasks such as passage retrieval, response generation, and citation labeling. RAD-Bench provides a frame to assess multi-turn LLMs’ capabilities in augmented generation with retrieved context in multi-turn scenarios with both Retrieval Synthesis and Retrieval Reasoning mechanisms. The evaluation framework contains 89 multi-turn question samples sampled across six practical scenarios inspired by human-LLM multi-turn dialogue interactions requiring retrieved context to complete tasks.